1. You deposit $1000 in an account today. You'll deposit $600 at the end of each month for the next 12 months & $800 each month for the following 12 months. How much interest will you have earned in 2 years if the account pays 5.5% compounded monthly?

(Answer is 962.57 but I'm not sure how to get there. Need both formula & financial calculator methods)

2. Two years ago you bought a bond with 2 years to maturity, face value $1000, semi-annual coupons, coupon rate 5%, YTM 2.5%. The Government of Canada is retiring this bond today, when the YTM on 2-year bonds equals 4%. What cash payment will you receive on the day the bond matures?

(answer is 1,025)

3. Jamie owes $21,750 at 5% rate of interest. The minimum amount she must pay monthly is $230.69. How much faster can she pay off this loan if she makes monthly payments of $300?

(answer is 2.79 years sooner)
Answer #1

1.

FORMULA:

=(1000*(1+5.5%/12)^24+600/(5.5%/12)*(1-1/(1+5.5%/12)^12)*(1+5.5%/12)^24+800/(5.5%/12)*(1-1/(1+5.5%/12)^12)*(1+5.5%/12)^12)-(1000+600*12+800*12)

=962.57

FINANCIAL CALCULATOR:

PV=1000

PMT=600

I/Y=5.5%/12

N=12

CPT FV=8440.71

Now

PV=8440.71

I/Y=5.5%/12

PMT=800

N=12

CPT FV=18762.57

Interest earned=18762.57-1000-600*12-800*12=962.57

2.

=1000+1000*5%/2=1025

3.

=NPER(5%/12,-230.69,21750)/12-NPER(5%/12,-300,21750)/12

=2.792053576